OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cos(pi/3)cos(pi/5) + sin(pi/3)sin(pi/5)

OpenStudy (cwrw238):

use the compound identity cos (a - b) = cosacosb +sinasinb

OpenStudy (cwrw238):

now insert a = pi/3 and b = pi/5 and work out lefthand side and you've got it

OpenStudy (cwrw238):

do you gollow me?

OpenStudy (anonymous):

kind of, not entirely

OpenStudy (cwrw238):

lol * follow i meant!

OpenStudy (anonymous):

haha i got what you meant

OpenStudy (cwrw238):

you just need to simplify cos (pi3 - pi/5)

OpenStudy (anonymous):

you need to get a common denominator right?

OpenStudy (cwrw238):